# Quantum (quadratic) gravity: replacing the massive tensor ghost with an inverted harmonic oscillator-like instability

### 07 Mar 2026 v1

@@ 1 Introduction
In nutshell, the present work resolves the longstanding tension between perturbative renormalizability and unitarity in quadratic gravity that arises from an incomplete understanding of the dynamical structure of the additional spin–2 sector. We show that for the physically consistent sign of the Weyl–squared coefficient, the extra tensor mode is not a ghost excitation but is governed by the Hamiltonian of a IHO-like degree of freedom (which we named dual IHO). Such systems describe controlled dynamical instabilities rather than negative–norm degrees of freedom, and they admit a consistent quantum description when formulated within a direct–sum Hilbert space structure. Within this framework, the spin–2 sector remains off–shell and contributes only through its dispersive, virtual effects, thereby preserving the renormalizability of quadratic gravity while avoiding violations of the optical theorem. The resulting theory retains the massless graviton and the scalaron as the only asymptotic propagating degrees of freedom, while the dual–IHO sector plays the role of a virtual UV completion.
@@
This paper is organized as follows. In Sec. 2 we discuss the HO, IHO and dual IHO physics and their role in QFT, SM, QFTCS. We highlight how quantum mechanics and QFT usually rely on the presumption of the arrow of time, but the occurrence of IHO instabilities challenges this view. In Sec. 3, we review the formulation of direct-sum quantum (field) theory which splits the quantum states as direct-sum of two components with two arrows of time in the parity conjugate geometric superselection sectors. We also detail on how this approach unveils a new understanding of quantum IHO and its connection of the spectrum with non-trivial zeros of the Riemann zeta function. In Sec. 4 we present a detailed review of quadratic gravity and its spin-0 and spin-2 degrees of freedom. We especially discuss the cases of spin-2 ghost and the dual IHO spin-2 cases exclusively. We discuss on the literature so-far on saving unitarity of quadratic gravity with spin-2 ghost and make heuristic comparisons between various approaches. In Sec. 5 we explore the case of dual IHO spin–2 field naturally restoring unitarity in quadratic gravity. We present the analysis of scattering amplitudes, 1-loop effects and renormalizability with the virtual contributions from dual IHO spin–2 field in the physical processes. We explicitly prove how the dual spin–2 field does not enter the unitarity cuts, does not go on-shell and consequently the optical theorem can be respected. We discuss the quantization of dual IHO spin–2 sector with DQFT approach, which cannot be interpreted as a particle and the asymptotic state. In Sec. 6, we discuss the finite action principle and its role in avoiding singular solutions in the unitary quadratic gravity. We argue that Starobinsky inflationary scenario could be an emergent in QQG since the dual IHO spin-2 sector can potentially drive the Universe towards a safe beginning. We analyze the UV behavior of 1-loop beta functions in unitary QQG and find the analogies with quantum electrodynamics (QED). In Sec. 7 we compute predictions of inflationary scenario in the unitary QQG which can be tested with future gravitational wave data. We derive, in particular, parity asymmetry of primordial power spectra that naturally emerges with DSI in QQG. In Sec. 8, we provide the highlights of our investigation with future outlook. The Appendix. A derives the effective action for the massless graviton during inflation.
@@
Figure 1: Phase space of the IHO (left panel) and the dual IHO (right panel) representing doubly degenerate positive and negative energy solutions in (2.16). The negative energy trajectories are given by Q>0,P<0Q>0,\,P<0 and Q<0,P>0Q<0,\,P>0 whereas the positive energy trajectories are Q>0,P>0Q>0,\,P>0 and Q<0,P<0Q<0,\,P<0. These double degenerate trajectories are related by (2.14) whereas the positive and negative energy regions are related by (2.15).
It is worth noticing that the IHO phase space can also be mimicked by another equivalent physical system which we call the dual IHO, whose Hamiltonian is given by
@@
The dual IHO Hamiltonian (2.12) is indefinite similar to the IHO Hamiltonian (2.5). The phase space of the dual IHO can be understood as a $90^{\circ}$ (clockwise) rotated version of IHO (See Fig. 1), where in each region the role of position and the momentum gets swapped (i.e., $E\to-E$) since the classical solutions of dual IHO are (See the right panel of Fig. 1)
@@
We witness an important symmetry here in both (2.10) and (2.13), the $\mathcal{P}\mathcal{T}$ (Parity $\mathcal{P}$ and Time reversal $\mathcal{T}$), i.e., $t\to-t,\,\tilde{q}\to-\tilde{q}$. The phase space of IHO and dual IHO are depicted in the left and right panels of Fig. 1 respectively. The regions $\rm I(III)$} and $\rm II(IV)$} are $\mathcal{P}\mathcal{T}$ conjugates of each other. The lines of $\tilde{p}=\pm\tilde{q}$ are called separatrices, which divide the classical trajectories into four distinct domains. We can notice that these regions come as degenerate pairs with opposite arrows of time with $\mathcal{P}\mathcal{T}$ symmetry
@@
The positive and negative energy regions of IHO and dual IHO in Fig. 1 are related by the following transformations
@@
The position- and momentum-space wavefunctions of the IHO Hamiltonian (which are equivalent for dual IHO, too) (for $Q>0$ and $E<0$, in units $\omega=1$) are given by [126, 29, 136]
@@
System Hamiltonian Nature
Harmonic oscillator $\frac{1}{2}\left(\tilde{p}^{2}+\tilde{q}^{2}\right)$ Positive definite
IHO $\frac{1}{2}\left(\tilde{p}^{2}-\tilde{q}^{2}\right)$ Indefinite
Ghost $\frac{1}{2}\left(\tilde{p}^{2}+\tilde{q}^{2}\right)$ Negative definite
Dual IHO $\frac{1}{2}\left(-\tilde{p}^{2}+\tilde{q}^{2}\right)$ Indefinite
Table 1: This table juxtaposes the nature of Hamiltonians of HO, IHO, ghost, and the dual IHO. We can notice that the IHO and dual IHO are identical physical systems which share the equivalent phase space (See Fig. 1).
@@ 2.1 Arrow of time in standard quantum theory and the challenge from IHO
In QM, Schrödinger equation is first order differential equation in time
@@
The IHO phase spaces in Fig. 1 pose us a new challenge where parity conjugate worlds split into the domains with opposite arrow of time. It has been an open problem since Berry and Keating [29] first looked at quantization of this physical system which poses a challenge to the standard Shrödinger equation’s practice of fixing an arrow of time in the first place. Berry and Keating suggested an antipodal identification $\left(Q,\,P\right)\leftrightarrow\left(-Q,\,-P\right)$ so that one can achieve periodicity with the $\mathcal{P}\mathcal{T}$ symmetry hyperbolic trajectories. This proposal in an ad-hoc way justified matching of IThis proposal, in an ad-hoc way, justified the matching of the IHO spectrum with the non-trivial zeros of the Riemann zeta function; the physical understanding of this proposal was an open question until our recent study [65]. harmonic oscillator (IHO) physics underlies the Higgs sector through the tachyonic instability, practical Standard Model (SM) calculations only involve fluctuations of the Higgs field around the true vacuum, where ordinary harmonic-oscillator (HO) physics and a particle interpretation apply. For this reason, the physical Higgs boson is understood as a fluctuation about the minimum of the potential. To explicitly see this, let us write down the Higgs potential
@@ 2.2 Occurrence of IHO physics in QFT in curved spacetime
As demonstrated in [65] quantum fields in BH spacetime has close correspondence with IHO physics. The exponentially growing and decaying character of $\left(U,,V\right)$ (2.37) generate the dilatation symmetries akin to the phase space of IHO we discussed in the last section (e.g., compare (2.16) and (2.37)). It is explicitly proved [33, 32, 136] that the wave function of a single state at the BH Horizon is characterized by the Berry and Keating type of IHO Hamiltonian (2.11).
@@ 2.3 Inflationary (quasi-de Sitter) quantum fluctuations
Inflationary quantum fluctuations also contain IHO physics. This is an interesting example where gravity is dynamical, and we can see how gravitational fluctuations propagate through the exponentially expanding spacetime and leave their imprints as temperature fluctuations in the CMB and gravitational waves. Inflation is approximately quasi-de Sitter [128], with mild departures encoded in the slow-roll parameters
@@ 3 The Direct-Sum Quantum (Field) Theory
and the propagator follows analogously. Interactions also split block-diagonally, e.g.
@@
so there is no mixing between sectors. As a result, all standard QFT amplitudes are reproduced in Minkowski space as a average of identical $\mathcal{P}\mathcal{T}$ symmetric quantities. In the case of a scattering process,
$A_{N\to M}=\frac{A^{N\to M}_{+}(p_{a},-p_{b})+A^{N\to M}_{-}(-p_{a},p_{b})}{2},\qquad A^{N\to M}_{+}=A^{N\to M}_{-},$ (3.19)
@@
3.1 Quantization of IHO or dual IHO in direct-sum quantum theory
Let us now define the new quantum construction of the IHO and dual IHO using a direct-sum framework [65], consistent with the key insights of Berry and Keating and Aneva [29, 9]. The IHO and dual IHO phase spaces (See Fig. 1) can be organized into regions related by a discrete set of transformations that form the dihedral group $D_{4}$ of order eight [9],
@@ 3.1 Quantization of IHO or dual IHO in direct-sum quantum theory
which generate the Riemann zeros $\zeta(\tfrac{1}{2}\pm i\bar{T})$. Although formally similar to the Berry–Keating condition (2.27), (3.30) now admits a geometric interpretation via superselection sectors, a key element missing in the original proposal [29, 9]. In the context of dual IHO, we must apply $E\to-E$ for the components of all regions in (3.25)-(3.30).
In summary, direct-sum quantization resolves the quantum chaos pathology of the IHO (or dual IHO) [31, 29] by geometrically splitting phase space into superselection sectors and connecting them through mathematical bridges (direct-sums) between sheets with opposite arrows of time. This yields a physically transparent realization of the Berry and Keating program and embeds the Riemann zero structure into a consistent quantum framework. The concept of mathematical bridges here echoes very well with the Einstein-Rosen bridges which is discussed in detail in [65].
@@ 4 Stelle or quadratic gravity and Renormalizability
Quadratic curvature modification of Einstein’s gravity is well known to improve the ultraviolet behavior of the theory and to render the gravitational action perturbatively renormalizable. However, when linearized around maximally symmetric backgrounds, such theories generically exhibit additional propagating degrees of freedom beyond the massless graviton. Among these, the appearance of a massive spin–2 mode with a tachyonic or negative mass-squared term has traditionally been interpreted as signaling a fundamental instability of the theory [44]. This interpretation has motivated extensive efforts to eliminate or evade such modes through parameter tuning, nonlocal modifications [^4] The non-local modifications of quadratic gravity are aimed to resolve the ghost problem by the choice of infinite derivative form factors which can be fine tuned further to achieve superrenormalizability [91, 134, 109, 35]. However, the non-local theories equations of motion are overly complicated which is a huge price to pay [90, 89]. Furthermore, recent investigations have indicated the need for selection of formfactors with background dependence to avoid ghosts [87, 88]. , or alternative quantization prescriptions. In this section, we revisit the physical meaning of the spin–2 tachyon in quadratic gravity and argue that its presence does not necessarily indicate a pathological instability. Instead, we propose that this mode admits a consistent reinterpretation as an inverted harmonic oscillator degree of freedom, naturally associated with spacelike momentum support and horizon-related quantum dynamics. Thus, the IHO-like spin–2 instability is rather healthier in contrast with pathological spin–2 ghost excitation case.

### 25 Mar 2026 v2

@@ 1 Introduction
In nutshell, the present work resolves the longstanding tension between perturbative renormalizability and unitarity in quadratic gravity that arises from an incomplete understanding of the dynamical structure of the additional spin–2 sector. We show that for the physically consistent sign of the Weyl–squared coefficient, the extra tensor mode is not a ghost excitation but is governed by the Hamiltonian of a IHO-like degree of freedom (which we named dual-IHO). Such systems describe controlled dynamical instabilities rather than negative–norm degrees of freedom, and they admit a consistent quantum description when formulated within a direct–sum Hilbert space structure. Within this framework, the spin–2 sector remains off–shell and contributes only through its dispersive, virtual effects, thereby preserving the renormalizability of quadratic gravity while avoiding violations of the optical theorem. The resulting theory retains the massless graviton and the scalaron as the only asymptotic propagating degrees of freedom, while the dual–IHO sector plays the role of a virtual UV completion.
@@
This paper is organized as follows. In Sec. 2 we discuss the HO, IHO and dual-IHO physics and their role in QFT, SM, QFTCS. We highlight how quantum mechanics and QFT usually rely on the presumption of the arrow of time, but the occurrence of IHO instabilities challenges this view. In Sec. 3, we review the formulation of direct-sum quantum (field) theory which splits the quantum states as direct-sum of two components with two arrows of time in the parity conjugate geometric superselection sectors. We also detail on how this approach unveils a new understanding of quantum IHO and its connection of the spectrum with non-trivial zeros of the Riemann zeta function. In Sec. 4 we present a detailed review of quadratic gravity and its spin-0 and spin-2 degrees of freedom. We especially discuss the cases of spin-2 ghost and the dual-IHO spin-2 cases exclusively. We discuss on the literature so-far on saving unitarity of quadratic gravity with spin-2 ghost and make heuristic comparisons between various approaches. In Sec. 5 we explore the case of dual-IHO spin–2 field naturally restoring unitarity in quadratic gravity. We present the analysis of scattering amplitudes, 1-loop effects and renormalizability with the virtual contributions from dual-IHO spin–2 field in the physical processes. We explicitly prove how the dual spin–2 field does not enter the unitarity cuts, does not go on-shell and consequently the optical theorem can be respected. We discuss the quantization of dual-IHO spin–2 sector with DQFT approach, which cannot be interpreted as a particle and the asymptotic state. In Sec. 6, we discuss the finite action principle and its role in avoiding singular solutions in the unitary quadratic gravity. We argue that Starobinsky inflationary scenario could be an emergent in QQG since the dual-IHO spin-2 sector can potentially drive the Universe towards a safe beginning. We analyze the UV behavior of 1-loop beta functions in unitary QQG and find the analogies with quantum electrodynamics (QED). In Sec. 7 we compute predictions of inflationary scenario in the unitary QQG which can be tested with future gravitational wave data. We derive, in particular, parity asymmetry of primordial power spectra that naturally emerges with DSI in QQG. In Sec. 8, we provide the highlights of our investigation with future outlook. The Appendix. A derives the effective action for the massless graviton during inflation.
@@
Figure 1: Phase space of the IHO (left panel) and the dual-IHO (right panel) representing doubly degenerate positive and negative energy solutions in (2.16). The negative energy trajectories are given by Q>0,P<0Q>0,\,P<0 and Q<0,P>0Q<0,\,P>0 whereas the positive energy trajectories are Q>0,P>0Q>0,\,P>0 and Q<0,P<0Q<0,\,P<0. These double degenerate trajectories are related by (2.14) whereas the positive and negative energy regions are related by (2.15).
It is worth noticing that the IHO phase space can also be mimicked by another equivalent physical system which we call the dual-IHO, whose Hamiltonian is given by
@@
The dual-IHO Hamiltonian (2.12) is indefinite similar to the IHO Hamiltonian (2.5). The phase space of the dual-IHO can be understood as a $90^{\circ}$ (clockwise) rotated version of IHO (See Fig. 1), where in each region the role of position and the momentum gets swapped (i.e., $E\to-E$) since the classical solutions of dual-IHO are (See the right panel of Fig. 1)
@@
We witness an important symmetry here in both (2.10) and (2.13), the $\mathcal{P}\mathcal{T}$ (Parity $\mathcal{P}$ and Time reversal $\mathcal{T}$), i.e., $t\to-t,\,\tilde{q}\to-\tilde{q}$. The phase space of IHO and dual-IHO are depicted in the left and right panels of Fig. 1 respectively. The regions $\rm I(III)$} and $\rm II(IV)$} are $\mathcal{P}\mathcal{T}$ conjugates of each other. The lines of $\tilde{p}=\pm\tilde{q}$ are called separatrices, which divide the classical trajectories into four distinct domains. We can notice that these regions come as degenerate pairs with opposite arrows of time with $\mathcal{P}\mathcal{T}$ symmetry
@@
The positive and negative energy regions of IHO and dual-IHO in Fig. 1 are related by the following transformations
@@
The position- and momentum-space wavefunctions of the IHO Hamiltonian (which are equivalent for dual-IHO, too) (for $Q>0$ and $E<0$, in units $\omega=1$) are given by [134, 29, 144]
@@
2.1 Normalizability of Eigenstates and existence of ground state in HO, Ghost, and (dual) IHO
Here we point out an important distinctions between HO, ghost and (dual) IHO physical systems with regard to the normalizability of the wave functions and the existence of ground state.
The Hamiltonian $H_{\rm HO}=\frac{\omega}{2}(\tilde{p}^{2}+\tilde{q}^{2})$ is bounded below, with discrete spectrum $E_{n}=\omega\!\left(n+\tfrac{1}{2}\right)$. The eigenstates are Hermite functions,
$\psi_{n}(\tilde{q})=\mathcal{N}_{n}\,H_{n}(\tilde{q})\,e^{-\tilde{q}^{2}/2},$ (2.18)
which belong to $L^{2}(\mathbb{R})$: they decay as $e^{-\tilde{q}^{2}/2}$ and are square-integrable. Every eigenstate, including the ground state $\psi_{0}=\pi^{-1/4}e^{-\tilde{q}^{2}/2}$, satisfies $\int|\psi_{n}|^{2}\,d\tilde{q}=1$. The spectrum is bounded below and the ground state is the unique normalizable state of lowest energy.
[…]
The IHO Hamiltonian $H_{\rm IHO}=\tfrac{\omega}{2}(\tilde{q}^{2}-\tilde{p}^{2})=\tfrac{\omega}{2}(\tilde{q}^{2}+\partial_{\tilde{q}}^{2})$ is fundamentally different from both the HO and the ghost. The eigenvalue equation reads
$\frac{\omega}{2}(\tilde{q}^{2}+\partial_{\tilde{q}}^{2})\,\psi_{E}=E\,\psi_{E}\qquad\Longrightarrow\qquad\partial_{\tilde{q}}^{2}\,\psi_{E}=\!\left(\tilde{q}^{2}-\frac{2E}{\omega}\right)\psi_{E}.$ (2.19)
[…]
The second term grows as $e^{+\tilde{q}^{2}/4}$ from -$\infty$ and is not square-integrable. Normalizability requires $\Gamma(-\nu)=0$, i.e. $\nu=0,1,2,\ldots$, which gives
$E=-i\omega\!\left(n+\tfrac{1}{2}\right),\qquad n=0,1,2,\ldots$ (2.21)
These energies are purely imaginary which means the IHO Hamiltonian must be non-Hermitian. If we demand the IHO Hamiltonian to be Hermitian (which we do in this work), then only real values of energies are allowed, for any real energies the IHO eigenmodes are non-normalizable. However, every eigenstate is only Dirac-delta normalizable: $\langle E|E^{\prime}\rangle=\delta(E-E^{\prime})$. This is categorically different from both the HO and the ghost. As we argued before and also in Fig. 1, since the dual-IHO is an identical physical system as IHO, the conclusions about ground state non-normalizability apply straightforwardly. We summarize all these concepts in Table. 1.
System Hamiltonian Nature Ground state in $L^{2}(\mathbb{R})$
Harmonic oscillator $\dfrac{1}{2}\!\left(\tilde{p}^{2}+\tilde{q}^{2}\right)$ Positive definite Yes
IHO $\dfrac{1}{2}\!\left(\tilde{p}^{2}-\tilde{q}^{2}\right)$ Indefinite No
Ghost $\dfrac{1}{2}\!\left(\tilde{p}^{2}+\tilde{q}^{2}\right)$ Negative definite Yes
dual-IHO $\dfrac{1}{2}\!\left(-\tilde{p}^{2}+\tilde{q}^{2}\right)$ Indefinite No
Table 1: Comparison of the HO, IHO, ghost, and dual-IHO Hamiltonians (in units $\omega=1$). The IHO and dual-IHO are identical physical systems sharing the same phase space (see Fig. 1).
@@ 2.2 Arrow of time in standard quantum theory and the challenge from IHO
In QM, Schrödinger equation is a first-order differential equation in time
@@
The IHO phase spaces in Fig. 1 pose us a new challenge where parity conjugate worlds split into the domains with opposite arrow of time. It has been an open problem since Berry and Keating [29] first looked at quantization of this physical system which poses a challenge to the standard Shrödinger equation’s practice of fixing an arrow of time in the first place. Berry and Keating suggested an antipodal identification $\left(Q,\,P\right)\leftrightarrow\left(-Q,\,-P\right)$ so that one can achieve periodicity with the $\mathcal{P}\mathcal{T}$ symmetry hyperbolic trajectories. This proposal, in an ad-hoc way, justified the matching of the IHO spectrum with the non-trivial zeros of the Riemann zeta function; the physical understanding of this proposal was an open question until our recent study [67]. harmonic oscillator (IHO) physics underlies the Higgs sector through the tachyonic instability, practical Standard Model (SM) calculations only involve fluctuations of the Higgs field around the true vacuum, where ordinary harmonic-oscillator (HO) physics and a particle interpretation apply. For this reason, the physical Higgs boson is understood as a fluctuation about the minimum of the potential. To explicitly see this, let us write down the Higgs potential
@@ 2.3 Occurrence of IHO physics in QFT in curved spacetime
As demonstrated in [67] quantum fields in BH spacetime has close correspondence with IHO physics. The exponentially growing and decaying character of $\left(U,V\right)$ (2.41) generate the dilatation symmetries akin to the phase space of IHO we discussed in the last section (e.g., compare (2.16) and (2.41)). It is explicitly proved [33, 32, 144] that the wave function of a single state at the BH Horizon is characterized by the Berry and Keating type of IHO Hamiltonian (2.11).
@@ 2.4 Inflationary (quasi-de Sitter) quantum fluctuations
Inflationary quantum fluctuations also contain IHO physics. This is an interesting example where gravity is dynamical, and we can see how gravitational fluctuations propagate through the exponentially expanding spacetime and leave their imprints as temperature fluctuations in the CMB and gravitational waves. Inflation is approximately quasi-de Sitter [136], with mild departures encoded in the slow-roll parameters (defined in the Einstein frame)
@@ 3 The Direct-Sum Quantum (Field) Theory
and the propagator, which is a time-ordered product of the fields $\phi_{\pm}$ in their respective vacua
$\Delta^{\pm}_{F}(x-y)$ $\equiv\langle 0_{\pm}|T\{\hat{\phi}_{\pm}\hat{\phi}_{\pm}\}|0_{\pm}\rangle$ (3.18)
$=\theta(x^{0}-y^{0})\,\hat{\phi}_{\pm}(\pm x)\hat{\phi}_{\pm}(\pm y)+\theta(y^{0}-x^{0})\,\hat{\phi}_{\pm}(\pm y)\hat{\phi}_{\pm}(\pm x)$
$=\int\frac{d^{4}x}{\left(2\pi\right)^{4}}\frac{\mp i}{k^{2}+m^{2}\mp i\epsilon}e^{ik\cdot\left(x-y\right)}$
In the DQFT, the fields $\hat{\phi}_{\pm}$ act on superselection sectors (corresponding to parity conjugate regions) with opposite temporal orientation. This ensures that correlation functions in the two sectors correspond to opposite arrows of time, while maintaining $\mathcal{PT}$ symmetry of the full theory.
The Feynman propagators in DQFT are defined sectorially, and there is no meaningful way to define a global Feynman propagator because parity conjugate regions share an opposite arrow of time. We can split the Feynman propagator in the following way
$\Delta^{\pm}_{F}(k)=\frac{\mp i}{k^{2}+m^{2}\mp i\epsilon}=\mp i{\rm PV}\left(\frac{1}{k^{2}+m^{2}}\right)+\pi\delta\left(k^{2}+m^{2}\right)$ (3.19)
using the Sokhotski–Plemelj–Fox identity
$\frac{1}{X\pm i\epsilon}=\mathrm{PV}\!\left(\frac{1}{X}\right)\mp i\pi\,\delta(X)$ (3.20)
where $\rm PV$ here stands for the Cauchy principal value.
Interactions in the DQFT picture split block-diagonally, e.g.
@@
so there is no mixing between sectors. As a result, all standard QFT amplitudes are reproduced in Minkowski space as an average of identical $\mathcal{P}\mathcal{T}$ symmetric quantities. In the case of a scattering process, observables are the squares of amplitudes integrated over the phase space, giving cross-sections, decay rates, and transition probabilities [44, 135]. The total amplitude square in the DQFT is the average of quantities evaluated in the parity conjugate superselection sectors
$|A_{N\to M}|^{2}=\frac{|A^{N\to M}_{+}(p_{a},-p_{b})|^{2}+|A^{N\to M}_{-}(-p_{a},p_{b})|^{2}}{2},\qquad|A^{N\to M}_{+}|^{2}=|A^{N\to M}_{-}|^{2},$ (3.22)
@@
3.1 Quantization of IHO or dual-IHO in direct-sum quantum theory
Let us now define the new quantum construction of the IHO and dual-IHO using a direct-sum framework [67], consistent with the key insights of Berry and Keating and Aneva [29, 9]. The IHO and dual-IHO phase spaces (See Fig. 1) can be organized into regions related by a discrete set of transformations that form the dihedral group $D_{4}$ of order eight [9],
@@ 3.1 Quantization of IHO or dual-IHO in direct-sum quantum theory
which generate the Riemann zeros $\zeta(\tfrac{1}{2}\pm i\bar{T})$. Although formally similar to the Berry–Keating condition (2.31), (3.33) now admits a geometric interpretation via superselection sectors, a key element missing in the original proposal [29, 9]. In the context of dual-IHO, we must apply $E\to-E$ for the components of all regions in (3.28)-(3.33).
In summary, direct-sum quantization resolves the quantum chaos pathology of the IHO (or dual-IHO) [31, 29] by geometrically splitting phase space into superselection sectors and connecting them through mathematical bridges (direct-sums) between sheets with opposite arrows of time. This yields a physically transparent realization of the Berry and Keating program and embeds the Riemann zero structure into a consistent quantum framework. The concept of mathematical bridges here echoes very well with the Einstein-Rosen bridges, which are discussed in detail in [67].
@@ 4 Stelle or quadratic gravity and Renormalizability
Quadratic curvature modification of Einstein’s gravity is well known to improve the ultraviolet behavior of the theory and to render the gravitational action perturbatively renormalizable. However, when linearized around maximally symmetric backgrounds, such theories generically exhibit additional propagating degrees of freedom beyond the massless graviton. Among these, the appearance of a massive spin–2 mode with a tachyonic or negative mass-squared term has traditionally been interpreted as signaling a fundamental instability of the theory [46]. This interpretation has motivated extensive efforts to eliminate or evade such modes through parameter tuning, nonlocal modifications [^4] The non-local modifications of quadratic gravity are aimed to resolve the ghost problem by the choice of infinite derivative form factors which can be fine-tuned further to achieve superrenormalizability [95, 142, 116, 36]. However, the non-local theories equations of motion are overly complicated which is a huge price to pay [94, 93]. Furthermore, recent investigations have indicated the need for selection of formfactors with background dependence to avoid ghosts [91, 92]. , or alternative quantization prescriptions. In this section, we revisit the physical meaning of the spin–2 tachyon in quadratic gravity and argue that its presence does not necessarily indicate a pathological instability. Instead, we propose that this mode admits a consistent reinterpretation as an inverted harmonic oscillator degree of freedom, naturally associated with spacelike momentum support and horizon-related quantum dynamics. Thus, the IHO-like spin–2 instability is rather healthier in contrast with pathological spin–2 ghost excitation case.